1. The newly constructed bridge in your town is a steel arch bridge 600 m in length How much does the total length of the roadway decking change between temperature extremes of 30°C and 50°C?


1
Expert's answer
2021-02-25T11:22:52-0500

We can find the change in length of the bridge from the formula:


"\dfrac{\Delta L}{L_0}=\alpha\Delta T,""\Delta L=\alpha L_0\Delta T,""\Delta L=11\cdot10^{-6}\ ^{\circ}C^{-1}\cdot600\ m\cdot(50^{\circ}C-30^{\circ}C)=0.132\ m=13.2\ cm."